# Learning Equilibria from Data: Provably Efficient Multi-Agent Imitation Learning

### v1

@@ 3.1 Necessity of $\mathcal{C}(\mu,\nu)$ in non-interactive MAIL
This theorem illustrates a fundamental limitation of BC in zero-sum Markov games. Specifically, it reveals that even perfect recovery of the Nash expert’s state visitation distribution, along with complete knowledge of the transition model, is insufficient for minimizing the Nash gap. The key insight is that a Nash equilibrium only guarantees robustness against unilateral deviations. As a result, regions of the Markov game that require joint deviations to be visited may remain underexplored by the expert, leaving the learner vulnerable in those regions. This can be seen in Fig. 1, if the learner has a (jointly) inaccurate policy in state $s_{1}$, the best response of the agents can change the expert path to exploit the opponent in the red path of the Markov Game and the the green one respectively. Notably, this phenomenon persists even when the transition model is known. This can be seen as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{\mathrm{xplt1}}},{\color[rgb]{0.0,0.51,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.51,0.5}S_{\mathrm{xplt2}}}$ and $S_{\mathrm{copy}}$ are sets of states, and each action combination leads to a different unique state, i.e. $\left|{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{\mathrm{xplt1}}}\cup{\color[rgb]{0.0,0.51,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.51,0.5}S_{\mathrm{xplt2}}}\cup S_{\mathrm{copy}}}\right|=\left|{{\mathcal{A}}}\right|\left|{{\mathcal{B}}}\right|$. This highlights the necessity of interactive Imitation Learning to explore strategically important but unobserved regions of the state space.
@@
Since the former assumption is rarely realistic, we later propose an interactive algorithm (Algorithm 2) that actively queries expert demonstrations to reduce the Nash gap of the resulting policy.
@@ 4 Avoiding the single deviation concentrability in interactive MAIL
Both of our algorithms address key limitations of the approaches proposed in BLADES and MALICE by Tang et al. (2024), as their methods require exponential compute and focus solely on error propagation analysis without providing convergence guarantees for the resulting policies. In contrast, our algorithms are accompanied by both convergence guarantees and polynomial computational cost. To motivate these algorithms, let us briefly revisit the structure of the original proof. In offline BC, we lack data corresponding to the best responses against the estimated expert policies. As a result, it is not feasible to directly estimate the expectation in Eq. 2. To circumvent this, we apply a change of measure at the cost of introducing the single deviation concentrability term.
Our first approach to overcome this limitation is to introduce a Best Response oracle, which enables sampling from the distributions $(\mu,\nu^{\operatorname{E}})$ and $(\mu^{\operatorname{E}},\nu)$, where $\mu\in\mathrm{br}(\widehat{\nu})$ and $\nu\in\mathrm{br}(\widehat{\mu})$, thereby allowing us to estimate the relevant expectations without incurring the concentrability coefficient. Formally, we have the following definition, also used in previous works (see e.g. Hellerstein et al. (2019)).
@@
Let $(\widehat{\mu},\widehat{\nu})$ be a pair of policies for a Markov Game ${\mathcal{G}}$. Then, a Best Response Oracle generates policies $\mu\in\mathrm{br}(\widehat{\nu})$ and $\nu\in\mathrm{br}(\widehat{\mu})$.
@@
Even under the assumption of being able to generate samples from $\mu^{\operatorname{E}},\nu$, where $\nu\in\mathrm{br}(\widehat{\mu})$, two problems remain. First of all, the optimization problem Eq. 3 is non-convex in $\widehat{\mu}$. Secondly, in order to estimate the minimizer $\widehat{\mu}$, we need to collect data from the occupancy measure of the policy pair $\mu^{\operatorname{E}},\nu$ for $\nu\in\mathrm{br}\left({\widehat{\mu}}\right)$, which depends on the minimizer itself.
@@
$\frac{1}{K}\sum_{k=1}^{K}\left\langle{d_{0}},{V^{\mu_{E},\nu^{\star}_{k}}-V^{\mu_{k},\nu^{\star}_{k}}}\right\rangle\leq\sqrt{\frac{\left|{{\mathcal{A}}_{max}}\right|}{K(1-\gamma)^{2}}\sum_{k=1}^{K}\mathbb{E}_{s\sim d^{\mu_{k},\nu^{*}_{k}}}\left[{\left\|{\mu_{E}(\cdot\mid s)-\mu_{k}(\cdot\mid s)}\right\|^{2}}\right]},$
@@ 4.1 Efficient algorithm with a best response oracle
       Query BR oracle to obtain $\mu_{k}^{*}\in\mathrm{br(\widehat{\nu}_{k})},\nu_{k}^{*}\in\mathrm{br(\widehat{\mu}_{k})}$ ;
       Sample $S^{\mu}_{k}\sim d^{\mu_{k},\nu_{k}^{*}}$, $A^{\mu}_{k}\sim\mu_{E}(\cdot\mid S^{\mu}_{k})$, $S^{\nu}_{k}\sim d^{\mu_{k}^{*},\nu_{k}}$, $A^{\nu}_{k}\sim\nu_{E}(\cdot\mid S^{\nu}_{k})$ ;
@@
With the best response oracle and the bound in Eq. 4 in place, we can aim at applying a no-regret algorithm to the loss sequence $\left\{{\mathbb{E}_{s\sim d^{\mu_{k},\nu^{\star}_{k}}}\left[{\left\|{\mu_{E}(\cdot\mid s)-\mu_{k}(\cdot\mid s)}\right\|^{2}}\right]}\right\}^{K}_{k=1}$. Since these losses are not directly observable, MAIL-BRO (see Algorithm 1) at each iteration performs a step of exponential weights updates with a stochastic unbiased gradient denoted by $g^{\mu}_{k}$ and $g^{\nu}_{k}$ for the two players respectively. These gradient estimates can be shown to have almost surely bounded noise too.
@@
Let us run Algorithm 1 for $K=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|^{2}\log\left|{{\mathcal{A}}_{\max}}\right|\log(1/\delta)}{(1-\gamma)^{4}\varepsilon^{4}}}\right)$ iterations with learning rate $\eta=\frac{2\left|{{\mathcal{S}}}\right|\log\left|{{\mathcal{A}}_{\max}}\right|}{K}$. Then, the sequence of policies $\left\{{\mu_{k},\nu_{k}}\right\}^{K}_{k=1}$ satisfies with probability at least $1-5\delta$ that $\frac{1}{K}\sum^{K}_{k=1}\max_{\mu\in\Pi}\left\langle{d_{0}},{V^{\mu,\nu_{k}}}\right\rangle-\min_{\nu\in\Pi}\left\langle{d_{0}},{V^{\mu_{k},\nu}}\right\rangle\leq\mathcal{O}(\varepsilon).$ Therefore, setting $\delta=\mathcal{O}(\epsilon)$ ensures that for a certain $\hat{k}\sim\mathrm{Unif}([K])$ it holds that $\mathbb{E}\left[{\mathrm{Nash}\text{-}\mathrm{Gap}(\mu_{\widehat{k}},\nu_{\widehat{k}})}\right]\leq\epsilon$. That is, $\mu_{\hat{k}},\nu_{\hat{k}}$ is an $\varepsilon$-Nash equilibrium in expectation.
The proof can be found in Appendix F. We observe that, compared to standard BC, the sample complexity now is of the order $\mathcal{O}(\epsilon^{-4})$, which is worse by a factor of $\epsilon^{-2}$. However, this trade-off allows us to completely avoid dependence on the single policy deviation concentrability coefficient in the MAIL-BRO upper bound. That is, MAIL-BRO is able to effectively recover an approximate equilibrium from expert data in a larger class of games.
@@ 4.2 Avoiding the best response oracle thanks to the maximum uncertainty response
Here, we introduce our algorithm MURMAIL (Algorithm 2) which can be applied in the most general setting where $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$ and the best response oracle is not available. The idea is again to start from (4), but instead of querying the Best Response Oracle, the objective is upper bounded by the maximum uncertainty policy. It is important to note that the exploration follows in a decentralized way, avoiding the curse of multi-agents by exploring induced MDPs instead of the original Markov Game.
@@
The following result states the theoretical guarantees for Algorithm 2. A proof can be found in Appendix F.
@@
Let us run Algorithm 2 for $K=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|^{2}\log\left|{{\mathcal{A}}_{\max}}\right|\log(1/\varepsilon)}{(1-\gamma)^{4}\varepsilon^{4}}}\right)$ outer iterations and $T=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|^{3}\left|{{\mathcal{A}}_{\max}}\right|^{3}\log(1/\varepsilon)}{(1-\gamma)^{8}\varepsilon^{4}}}\right)$ inner iterations with learning rate $\eta=\frac{2\left|{{\mathcal{S}}}\right|\log\left|{{\mathcal{A}}_{\max}}\right|}{K}$. Then, for a certain $\hat{k}\sim\mathrm{Unif}([K])$ it holds that $\mathbb{E}\left[{\mathrm{Nash}\text{-}\mathrm{Gap}(\mu_{\widehat{k}},\nu_{\widehat{k}})}\right]\leq\epsilon$.
@@
Note that our algorithms scale with $\mathrm{poly}(\left|{{\mathcal{A}}_{\max}}\right|)$. While this may appear suboptimal in the two-player zero-sum setting, it is important to emphasize that the underlying algorithms support decentralized execution. In particular, in Algorithm 1, the dependence on ${\mathcal{A}}_{\max}^{2}$ does not stem from the two-player structure, but rather from the reformulation of the objective necessary to obtain an unbiased estimator for the gradient update. Similarly, the $\left|{{\mathcal{A}}_{\max}}\right|^{5}$ dependence in Algorithm 2 arises from the squared objective and the RL inner loop. Crucially, in this inner loop, each agent solves a single-agent MDP, ensuring that the algorithm remains fully decentralized. Altogether, these observations indicate that our algorithms scale linearly with $\left|{{\mathcal{A}}_{\max}}\right|$ and do not suffer from the curse of multi-agents in the $n$-player setting. A sketched version for $n$-player general-sum games can be found in Appendix I.
@@ 5 Numerical Validation
In this section, we provide a numerical evaluation of our proposed algorithms in the Markov Game considered in the lower bound construction (Fig. 1) as this environment allows us to control ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ by considering different convex combinations of the two pure Nash equilibria profiles (i.e., the black and the blue path in Figure 1). This environment serves as a proof of concept to demonstrate the practical feasibility of our methods. In particular, we aim to highlight that the performance of BC depends on the concentrability coefficient ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ even when it is bounded, and completely fails when ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$.
We evaluate Multi-Agent BC and MURMAIL (Algorithm 2) in the considered environment and measure the exploitability of the resulting policies with respect to the number of expert queries (for MURMAIL) and dataset size (for BC). The results are presented in Fig. 2.
@@
As predicted by our theoretical analysis, Multi-Agent BC fails in settings with ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$, whereas MURMAIL still succeeds in minimizing the Nash gap. However, in environments where ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty$, BC can outperform MURMAIL in terms of efficiency $\epsilon^{-2}$ compared to $\epsilon^{-8}$. Nevertheless, one should also consider that the performance of MURMAIL is independent of ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ and therefore MURMAIL can outperform BC in cases where ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is bounded but large. This highlights the importance of algorithm selection based on the underlying environment. Additional details, experiments in another environment, and practical insights for improving MURMAIL’s performance are discussed in Appendix K.

### v2

@@ 3.1 Necessity of $\mathcal{C}(\mu_{E},\nu_{E})$ in non-interactive MAIL
This theorem illustrates a fundamental limitation of BC in zero-sum Markov games. Specifically, it reveals that even perfect recovery of the Nash expert’s state visitation distribution, along with complete knowledge of the transition model, is insufficient for minimizing the Nash gap. The key insight is that a Nash equilibrium only guarantees robustness against unilateral deviations. As a result, regions of the Markov game that require joint deviations to be visited may remain underexplored by the expert, leaving the learner vulnerable in those regions. This can be seen in Fig.˜1, if the learner has a (jointly) inaccurate policy in state $s_{1}$, the best response of the agents can change the expert path to exploit the opponent in the red path of the Markov Game and the green one respectively. Notably, this phenomenon persists even when the transition model is known. This can be seen as ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{\mathrm{xplt1}}},{\color[rgb]{0.0,0.51,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.51,0.5}S_{\mathrm{xplt2}}}$ and $S_{\mathrm{copy}}$ are sets of states, and each action combination leads to a different unique state, i.e. $\left|{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}S_{\mathrm{xplt1}}}\cup{\color[rgb]{0.0,0.51,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.51,0.5}S_{\mathrm{xplt2}}}\cup S_{\mathrm{copy}}}\right|=\left|{{\mathcal{A}}}\right|\left|{{\mathcal{B}}}\right|$. This highlights the necessity of interactive Imitation Learning to explore strategically important but unobserved regions of the state space.
@@
Since the former assumption is rarely realistic, we later propose an interactive algorithm (Algorithm˜2) that actively queries expert demonstrations to reduce the Nash gap of the resulting policy.
@@ 4 Avoiding the single deviation concentrability in interactive MAIL
Both of our algorithms address key limitations of the approaches proposed in BLADES and MALICE by Tang et al. (2024), as their methods require exponential compute and focus solely on error propagation analysis without providing convergence guarantees for the resulting policies. In contrast, our algorithms are accompanied by both convergence guarantees and polynomial computational cost. To motivate these algorithms, let us briefly revisit the structure of the original proof. In offline BC, we lack data corresponding to the best responses against the estimated expert policies. As a result, it is not feasible to directly estimate the expectation in Eq.˜2. To circumvent this, we apply a change of measure at the cost of introducing the all deviation concentrability term.
[…]
@@
[…]
@@
[…]
@@
$\frac{1}{K}\sum_{k=1}^{K}\left\langle{d_{0}},{V^{\mu_{E},\nu^{\star}_{k}}-V^{\mu_{k},\nu^{\star}_{k}}}\right\rangle\leq\sqrt{\frac{\left|{{\mathcal{A}}_{max}}\right|}{K(1-\gamma)^{2}}\sum_{k=1}^{K}\mathbb{E}_{s\sim d^{\mu_{k},\nu^{\star}_{k}}}\left[{\left\|{\mu_{E}(\cdot\mid s)-\mu_{k}(\cdot\mid s)}\right\|^{2}}\right]},$
@@ 4.1 Efficient algorithm with a best response oracle
   Query BR oracle to obtain $\mu_{k}^{\star}\in\mathrm{br(\widehat{\nu}_{k})},\nu_{k}^{\star}\in\mathrm{br(\widehat{\mu}_{k})}$ ;
   Sample $S^{\mu}_{k}\sim d^{\mu_{k},\nu_{k}^{\star}}$, $A^{\mu}_{k}\sim\mu_{E}(\cdot\mid S^{\mu}_{k})$, $S^{\nu}_{k}\sim d^{\mu_{k}^{\star},\nu_{k}}$, $A^{\nu}_{k}\sim\nu_{E}(\cdot\mid S^{\nu}_{k})$ ;
@@
With the best response oracle and the bound in Eq.˜4 in place, we can aim at applying a no-regret algorithm to the loss sequence $\left\{{\mathbb{E}_{s\sim d^{\mu_{k},\nu^{\star}_{k}}}\left[{\left\|{\mu_{E}(\cdot\mid s)-\mu_{k}(\cdot\mid s)}\right\|^{2}}\right]}\right\}^{K}_{k=1}$. Since these losses are not directly observable, MAIL-BRO (see Algorithm˜1) at each iteration performs a step of exponential weights updates with a stochastic unbiased gradient denoted by $g^{\mu}_{k}$ and $g^{\nu}_{k}$ for the two players respectively. These gradient estimates can be shown to have almost surely bounded noise too.
@@
Let us run Algorithm˜1 for $K=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|^{2}\log\left|{{\mathcal{A}}_{\max}}\right|\log(1/\delta)}{(1-\gamma)^{4}\varepsilon^{4}}}\right)$ iterations with learning rate $\eta=\frac{2\left|{{\mathcal{S}}}\right|\log\left|{{\mathcal{A}}_{\max}}\right|}{K}$. Then, the sequence of policies $\left\{{\mu_{k},\nu_{k}}\right\}^{K}_{k=1}$ satisfies with probability at least $1-5\delta$ that $\frac{1}{K}\sum^{K}_{k=1}\max_{\mu\in\Pi}\left\langle{d_{0}},{V^{\mu,\nu_{k}}}\right\rangle-\min_{\nu\in\Pi}\left\langle{d_{0}},{V^{\mu_{k},\nu}}\right\rangle\leq\mathcal{O}(\varepsilon).$ Therefore, setting $\delta=\mathcal{O}(\varepsilon)$ ensures that for a certain $\hat{k}\sim\mathrm{Unif}([K])$ it holds that $\mathbb{E}\left[{\mathrm{Nash-}\mathrm{Gap}(\mu_{\widehat{k}},\nu_{\widehat{k}})}\right]\leq\varepsilon$. That is, $\mu_{\hat{k}},\nu_{\hat{k}}$ is an $\varepsilon$-Nash equilibrium in expectation.
The proof can be found in Appendix˜E. We observe that, compared to standard BC, the sample complexity now is of the order $\mathcal{O}(\varepsilon^{-4})$, which is worse by a factor of $\varepsilon^{-2}$. However, this trade-off allows us to completely avoid dependence on the all policy deviation concentrability coefficient in the MAIL-BRO upper bound. That is, MAIL-BRO is able to effectively recover an approximate equilibrium from expert data in a larger class of games.
@@ 4.2 Avoiding the best response oracle thanks to the maximum uncertainty response
Here, we introduce our algorithm MURMAIL (Algorithm˜2) which can be applied in the most general setting where $\mathcal{C}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$ and the best response oracle is not available. The idea is again to start from (4), but instead of querying the Best Response Oracle, the objective is upper bounded by the maximum uncertainty policy. It is important to note that the exploration follows in a decentralized way, avoiding the curse of multi-agents by exploring induced MDPs instead of the original Markov Game.
@@
The following result states the theoretical guarantees for Algorithm˜2. A proof can be found in Appendix˜E.
@@
Let us run Algorithm˜2 for $K=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}_{\max}}\right|^{2}\log\left|{{\mathcal{A}}_{\max}}\right|\log(1/\varepsilon)}{(1-\gamma)^{4}\varepsilon^{4}}}\right)$ outer iterations and $T=\mathcal{O}\left({\frac{\left|{{\mathcal{S}}}\right|^{3}\left|{{\mathcal{A}}_{\max}}\right|^{3}\log(1/\varepsilon)}{(1-\gamma)^{8}\varepsilon^{4}}}\right)$ inner iterations with learning rate $\eta=\frac{2\left|{{\mathcal{S}}}\right|\log\left|{{\mathcal{A}}_{\max}}\right|}{K}$. Then, for a certain $\hat{k}\sim\mathrm{Unif}([K])$ it holds that $\mathbb{E}\left[{\mathrm{Nash-}\mathrm{Gap}(\mu_{\widehat{k}},\nu_{\widehat{k}})}\right]\leq\varepsilon$.
@@
Note that our algorithms scale with $\mathrm{poly}(\left|{{\mathcal{A}}_{\max}}\right|)$. While this may appear suboptimal in the two-player zero-sum setting, it is important to emphasize that the underlying algorithms support decentralized execution. In particular, in Algorithm˜1, the dependence on ${\mathcal{A}}_{\max}^{2}$ does not stem from the two-player structure, but rather from the reformulation of the objective necessary to obtain an unbiased estimator for the gradient update. Similarly, the $\left|{{\mathcal{A}}_{\max}}\right|^{5}$ dependence in Algorithm˜2 arises from the squared objective and the RL inner loop. Crucially, in this inner loop, each agent solves a single-agent MDP, ensuring that the algorithm remains fully decentralized. Altogether, these observations indicate that our algorithms scale linearly with $\left|{{\mathcal{A}}_{\max}}\right|$ and do not suffer from the curse of multi-agents in the $n$-player setting. A sketched version for $n$-player general-sum games can be found in Appendix˜H.
@@ 5 Numerical Validation
In this section, we provide a numerical evaluation of our proposed algorithms in the Markov Game considered in the lower bound construction (Fig.˜1) as this environment allows us to control ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ by considering different convex combinations of the two pure Nash equilibria profiles (i.e., the black and the blue path in Figure˜1). This environment serves as a proof of concept to demonstrate the practical feasibility of our methods. In particular, we aim to highlight that the performance of BC depends on the concentrability coefficient ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ even when it is bounded, and completely fails when ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$. Note that in all considered cases, we have that $\beta=0$ and therefore the BC bound proven in Tang et al. (2024) would always be vacuous, while Theorem˜3.1 remains valid. The code used for the experiments is available at https://github.com/tfreihaut/Murmail.
We evaluate Multi-Agent BC and MURMAIL (Algorithm˜2) in the considered environment and measure the exploitability of the resulting policies with respect to the number of expert queries (for MURMAIL) and dataset size (for BC). The results are presented in Fig.˜2.
@@
As predicted by our theoretical analysis, Multi-Agent BC fails in settings with ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})=\infty$, whereas MURMAIL still succeeds in minimizing the Nash gap. However, in environments where ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})<\infty$, BC can outperform MURMAIL in terms of efficiency $\varepsilon^{-2}$ compared to $\varepsilon^{-8}$. Nevertheless, one should also consider that the performance of MURMAIL is independent of ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ and therefore MURMAIL can outperform BC in cases where ${\mathcal{C}}(\mu^{\operatorname{E}},\nu^{\operatorname{E}})$ is bounded but large. This highlights the importance of algorithm selection based on the underlying environment. Additional details, experiments in another environment, and practical insights for improving MURMAIL’s performance are discussed in Appendix˜J.